OpenStudy (anonymous):

Which system of equations is represented by the following graph? a graph of two intersecting lines on a coordinate plane. One passes through points (0, 3) and (6, 0). The other passes through points (0, −3) and (3, 0). x − 2y = 6 x + y = 3 y = x + 3 y = −1 over 2x + 3 y = x − 3 y =1 over 2x + 3 x + 2y = 6 x − y = 3

OpenStudy (anonymous):

any idea ?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

find the equation of the lines 1- find the slope of the line slope(m)= (y2-y1)/(x2-x1) m=(0-3)/(6-0)=-3/6=-1/2

OpenStudy (anonymous):

2- find the equation y-y1=m(x-1) pick any points of the two y-0=-1/2(x-6) y=-1/2x+3 multiply both sides by 2 2y=-x+6 x+2y=6 this is the first equation

OpenStudy (anonymous):

the second equation m=(0-(-3))/(3-0)=3/3=1 y-y1=m(x-x1) using second point y-0=1(x-3) y=x-3 this is the second equation

OpenStudy (anonymous):

-x+y=-3 multiply both sides by -1 x-y=3

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

You can get it by moving backwards. Check the system that satisfy the points while replacing their coordinates :)